OpenStudy (ayubie):

In order to use L'Hospital's rule, what does the limit need to be equal to? in other words, do I need to get 0/0, or can I still use it if I get 1/0, and 0/1?

OpenStudy (anonymous):

It must be \(\dfrac{0}{0}\) or \(\pm\dfrac{\infty}{\infty}\).

OpenStudy (anonymous):

\(\dfrac{0}{1}\) is just 0, and \(\dfrac{1}{0}\) is undefined, or one of \(\infty\) or \(-\infty\) depending on the direction of approach.

OpenStudy (mathmale):

Why not choose and post a few example problems, so that we could discuss what to do in specific cases? Then perhaps the general rules about when to use l'Hopital's Rule may make more sense.